OpenStudy (anonymous):

FIND THE VALUE OF: sin 15 degrees In trigonometry ADDITION AND SUBTRACTION IDENTITIES

OpenStudy (anonymous):

.6502

OpenStudy (anonymous):

It's not like that

OpenStudy (anonymous):

Try using a unit circle

OpenStudy (mertsj):

\[\sin \frac{30}{2}=\sqrt{\frac{1-\cos 30}{2}}\]

OpenStudy (anonymous):

\[\sin(15)=\sin(45-30)=\sin(45)\cos(30)-\cos(45)(\sin(30)\]

OpenStudy (anonymous):

\[=\frac{ 1 }{ \sqrt{2} }*\frac{ \sqrt{3} }{ 2 }-\frac{ 1 }{ \sqrt{2} }*\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Which is which @Mertsj @mustafa2014

OpenStudy (ybarrap):

$$ \sin 15=\sin ( 45-30)=\sin(45)\cos(30)-\cos(45)\sin(30)\\ =\sqrt2/2\times(\sqrt3/2)-\sqrt2/2\times(1/2)\\ =\sqrt2/2\cfrac{\sqrt3-1}{2}\\ =\sqrt2\cfrac{\sqrt3-1}{4}\\ =\sqrt2\cfrac{\sqrt3-1}{2\times 2}\\ =\cfrac{\sqrt3-1}{\sqrt2\times 2}\\ $$

OpenStudy (anonymous):

Cant i use sin(60-45)???

OpenStudy (anonymous):

you can use it :)

OpenStudy (anonymous):

What other values could i use @mustafa2014